Caffeine has a therapeutic index of approximately 100:1. If the ED50 for caffeine is 12 mg/kg,...

  1. Caffeine has a therapeutic index of approximately 100:1. If the ED50 for caffeine is 12 mg/kg, what is the LD50?
  2. In a typical dose-response curve graph, what is placed along the vertical, y-axis?
  3. Sheila weighs 55 kg (121 pounds). She starts her day off with a large cup of Starbucks coffee (330 mg caffeine). Every day she drinks at least three large cans of Monster energy drink, each containing 160 mg of caffeine. In the evening, she drinks two cups of Lipton tea, with 55 mg of caffeine each. That's a total of 920 mg of caffeine per day! What is Emily's daily caffeine dose?
  4. Heroin has a half-life of approximate 10 minutes. Ronaldo injects 10mg of heroin IV. How much heroin will remain in his system 20 mins after injection?
  5. Drug A has an ED50 of 10mg/kg and an LD50 of 100 mg/kg what is the therapeutic index of this drug?
  6. What is the half-life of caffeine?
  7. Drug X has an ED50 of 2 mg/kg and an LD50 of 400mg/kg. What is the therapeutic index for drug X?  
  8. At the LD 50 of a drug what % of the subjects would die?
  9. What is the threshold on a dose response curve? (If hard to find in the book feel free to use the internet.)
  10. If you were prescribed an extended release stimulant medication for ADHD and you took it at 8 in the morning, at about what time would the effects likely wear off?
Answer #1

Thearapeutic index.

It is a qualitative measurement.It is the ratio which explains concentration where the drug becomes effective and it become toxic.

LD 50/ED 50.Here the thearapeutic index is given as 100:1.

So LD 50/12 =100/1.Therefore LD 50=1200 mg/kg.

2)Average response which is recorded is placed on vertical ,y-axis.

The range of doses is placed on x-axis.